OpenStudy (b77w):

How am I supposed to find the solution for this system? 4x+2y+z=10 -6x-y-5z=-34 3x+3y+6z=9

OpenStudy (amistre64):

by any of the assorted techniques that your studies have developed

OpenStudy (amistre64):

id start by reducing the rows

OpenStudy (amistre64):

4x+2y+z=10 -6x-y-5z=-34 3x+3y+6z=9 4x + 2y + z = 10 6x + y + 5z = 34 x + y + 2z = 3

OpenStudy (amistre64):

i see we can have a y solucion in terms of x and z by using the last 2

OpenStudy (amistre64):

6x + y + 5z = 34 x + y + 2z = 3 6x + 5z = 34 x + 2z = 3

OpenStudy (amistre64):

or forget the drama of substitution and go right for the matrix juggular rref{{4,2,1,10},{6,1,5,34},{1,1,2,3}} http://www.wolframalpha.com/input/?i=rref%7B%7B4%2C2%2C1%2C10%7D%2C%7B6%2C1%2C5%2C34%7D%2C%7B1%2C1%2C2%2C3%7D%7D

OpenStudy (amistre64):

the right column produces the results

OpenStudy (b77w):

I need to solve with substitution or elimination

OpenStudy (amistre64):

the matrix row reduction is defined by elementary row operations that mirror the sub/elim method

OpenStudy (amistre64):

solve 1 equation for one variable in terms of the other 2; sub that into the other 2 equation; resolve the issue with the 2 unkown 2 variable elimination and work that back into the 1st equation

OpenStudy (amistre64):

or define the line of intersection of the first 2 planes and then calculate where that would hit the third plane

OpenStudy (b77w):

I tried that. I keep getting weird numbers

OpenStudy (amistre64):

then your algebra skills need to be honed more :)

OpenStudy (b77w):

I just want to see someone else's steps. I know what the answers are

OpenStudy (amistre64):

4x + 2y + z = 10 ; z=10-4x-2y, sub it in 6x + y + 5(10-4x-2y) = 34 x + y + 2(10-4x-2y) = 3 6x + y + 50 -20x -10y = 34 x + y + 20 - 8x - 4y = 3 reduce those equations for me

OpenStudy (b77w):

what does that mean?

OpenStudy (anonymous):

elimination 4x + 2y + z = 10 -6x - y - 5z = -34 -->(2) --------------- 4x + 2y + z = 10 -12x - 2y - 10z = -68 (result of multiplying by 2) ---------------add -8x - 9z = -58 -6x - y - 5z = -34 -->(3) 3x + 3y + 6z = 9 --------------- -18x - 3y - 15z = - 102 (result of multiplying by -3) 3x + 3y + 6z = 9 ----------------add -15x - 9z = - 93 -8x - 9z = -58 -15x - 9z = -93 -->(-1) --------------- -8x - 9z = -58 15x + 9z = 93 --------------add 7x = 35 x = 5 -8x - 9z = -58 -8(5) - 9z = -58 -40 - 9z = -58 -9z = -58 + 40 -9z = - 18 z = 2 4x + 2y + z = 10 4(5) + 2y + 2 = 10 20 + 2y + 2 = 10 22 + 2y = 10 2y = 10 - 22 2y = -12 y = -6 check... 3x + 3y + 6z = 9 3(5) + 3(-6) + 6(2) = 9 15 - 18 + 12 = 9 27 - 18 = 9 9 = 9 (correct) x = 5, y = -6, z = 2 or (5,-6,2)

OpenStudy (anonymous):

do you have any questions ?